OpenStudy (ohmybookness):

In which of the three forms is this function written: y+8=2(x-1)^2

OpenStudy (ohmybookness):

The lesson is on Quadratic functions if this helps?

OpenStudy (anonymous):

It would be a vertex form. The three forms of quadratic functions I guess you're referring to are the standard, vertex and intercept forms where standard form looks like y = ax^2 + bx +c, vertex looks like y = a(x-1)^2 + b, intercept form looks like y = a(x-2)(x-4)

OpenStudy (ohmybookness):

Thank You! Do you think you can help me out with a couple more?

OpenStudy (ohmybookness):

It says that i need to convert it into standard form and to show my work. How would that look?

OpenStudy (anonymous):

To make it into standard form you would just expand the equation. So for this equation: y+8=2(x-1)^2 y = 2(x-1)^2 - 8 y = 2 * (x^2 - 2x + 1) - 8 y = 2x^2 - 4x + 2 - 8 y = 2x^2 - 4x - 6

OpenStudy (ohmybookness):

would the vertex of the parabola be (1/ -8)

OpenStudy (anonymous):

Yep, (1, -8) - move one up and 8 units left from (0,0)

OpenStudy (ohmybookness):

It says to use the formula for finding the axis of symmetry and vertex of a parabola in standard form. Does it match the vertex being (1,-8)

OpenStudy (anonymous):

Oops, sorry *one unit right, 8 units down

OpenStudy (anonymous):

Since vertex is at (1, -8), axis of symmetry would be x = 1

OpenStudy (anonymous):